OpenStudy (anonymous):

Solve the DE (2x+1)^2 y''+8(2x+1) y'+5y=0 Hint: let t=2x+1

OpenStudy (anonymous):

Another Euler-Cauchy equation. Assume a solution is of the form \(y=t^r\).

OpenStudy (anonymous):

I got the roots as \[\frac{ -7 \pm \sqrt{29}}{2 }\] Do i set that up using y(x)=c1 cos(ln x) + c2 sin (ln x)

OpenStudy (anonymous):

No, you don't have any complex roots, there's no need for Euler's formula here.

OpenStudy (anonymous):

Thanks again.

OpenStudy (anonymous):

yw